# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v2

@@ 7.1. Unique descending chains
Proof of (1): Let $N=\tilde{L}^{N_{G}(H)}$. Thus $L/N$ is Galois as $H\unlhd N_{G}(H)$. We have $[L:N]=[N_{G}(H):H]$. Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $H\unlhd{\rm Gal}(\tilde{L}/N^{\prime})\subseteq G$. Thus ${\rm Gal}(\tilde{L}/N^{\prime})\subseteq N_{G}(H)$. Hence $N\subseteq N^{\prime}$. Since $[L:N]=[L:N^{\prime}]$, we have $N=N^{\prime}$.
Proof of (2): Let $N_{0}=L$, From (1), we can inductively choose $N_{i}$ for each $i\geq 1$. Let $N_{i}/K$ be the unique intermediate extension of $N_{i-1}/K$ such that $N_{i}/N_{i-1}$ is Galois with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$. The chain terminates since $L/K$ is finite.
@@
Alternate proof for Thm 7.1.1 (1).
Let $N=L^{{\rm Aut}(L/K)}$. Hence $L/N$ is Galois and $[L:N]=|{\rm Aut}(L/K)|$. Hence by Thm 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $N^{\prime}=L^{{\rm Aut}(L/N^{\prime})}$ and $[L:N^{\prime}]=|{\rm Aut}(L/N^{\prime})|$. Hence $|{\rm Aut}(L/N^{\prime})|=|{\rm Aut}(L/K)|$. By Prop 6.1.1 (1), ${\rm{\rm Aut}}(L/N^{\prime})\subset{\rm Aut}(L/K)$. Thus ${\rm{\rm Aut}}(L/N^{\prime})={\rm Aut}(L/K)$. Hence $N^{\prime}=N$.
@@
Equivalently we can state Thm 7.1.1 (1) as follows: There exists a unique intermediate extension $N/K$ such that $L/N$ is Galois of maximum possible degree. This is because, since $L/N$ is Galois, we have $[L:N]=|{\rm Aut}(L/N)|$ and we also have ${\rm{\rm Aut}}(L/N)\subset{\rm Aut}(L/K)$. Hence $[L:N]$ is bounded by $r_{K}(L)$.
@@
Example 7.1.4.
Let the notation and conditions be as in Example 5.1.3. Let $L=\mathbb{Q}(a)$ and $N_{i}=\mathbb{Q}(a^{2^{i}})$. We compute and show that the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{v_{2}(n)}$ where $v_{2}$ is the standard $2$-adic valuation.
If $n$ is odd then $r_{K}(L)=1$. Hence by Thm 7.1.1(5), the unique descending chain is singleton $L$.
If $n$ is even then $r_{K}(L)=2$. Identifying $G={\rm Gal}(\mathbb{Q}_{f}/\mathbb{Q})$ with $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ we have $H=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}\subseteq G$. Hence $N_{G}(H)=\mathbb{Z}/2\mathbb{Z}\times(\mathbb{Z}/n\mathbb{Z})^{\times}$ where $\mathbb{Z}/2\mathbb{Z}$ is generated by the element of $G$ mapping $a$ to $-a$ and $b$ to $b$. Hence $\tilde{L}^{N_{G}(H)}=\mathbb{Q}(a^{2})=N_{1}$.
Now $\tilde{N_{1}}=\mathbb{Q}(a^{2},b^{2})$. Also since $\mathbb{Q}(a)\cap\mathbb{Q}(b)=\mathbb{Q}$. Hence $\mathbb{Q}(a^{2})\cap\mathbb{Q}(b^{2})=\mathbb{Q}$. Hence ${\rm Gal}(\tilde{N_{1}}/K)$ can be identified with $\mathbb{Z}/(n/2)\mathbb{Z}\rtimes(\mathbb{Z}/(n/2)\mathbb{Z})^{\times}$. If $n/2$ is even, we can repeat the process for $N_{1}$ and get $N_{2}$. The process will terminate at $N_{v_{2}(n)}$ as $n/v_{2}(n)$ is odd.
Remark 7.1.5.
The above is also true for the case not covered, that is $n$ is even and $\sqrt{c}\in\mathbb{Q}(b)$. The following proof works in general. We know that $N_{1}=L^{{\rm Aut}(L/K)}$. Since $n$ is even, ${\rm{\rm Aut}}(L/K)$ has $2$ elements mapping $a$ to $a$ and $a$ to $-a$. Hence $\mathbb{Q}(a^{2})\subset L^{{\rm Aut}(L/K)}$. Now since $x^{n}-c$ is minimal polynomial for $a$ over $\mathbb{Q}$. Hence $a^{2}$ satisfies $x^{n/2}-c$. We claim that it is indeed the minimal polynomial for $a^{2}$ over $\mathbb{Q}$. If not, then let $f(x)$ be minimal polynomial for $a^{2}$ with degree $<n/2$. Then $a$ satisfies the polynomial $f(x^{2})$ which has degree $<n$ which gives a contradiction. Hence $[\mathbb{Q}(a^{2}):\mathbb{Q}]=n/2$. Thus $N_{1}=\mathbb{Q}(a^{2})$. Proceeding similarly we get the unique descending chain.
Example 7.1.6.
[…]
We now state a result that describes the unique descending chain for the extension considered in Thm. 3.2.4 more precisely.
Theorem 7.1.7.
Let $L_{k}$ be as in proof of Thm. 3.2.4. Let $N_{k}$ be the unique intermediate extension for $L_{k}/K$ as in Thm. 7.1.1 (1). Then we have the following.
  1. (1)
  2. (2)
    Case (a) Characteristic of $K$ $\neq 2$. For $k<n-1$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
    Case (b) Characteristic of $K=2$. For $k<n-1$ but $k\neq n/2$, we have $N_{k}=K(t_{1})=K(\alpha_{1}+\alpha_{2}\dots+\alpha_{k})$.
  3. (3)
    for $k<n/2$, the unique descending chain is $L_{k}\supsetneq N_{k}$.
  4. (4)
    for $k=n/2$ and for field K with characteristic $\neq 2$, the cluster size $r_{K}(N_{k})=2$.
@@
(1) Now, $\tilde{L_{k}}=L_{n-1}=L_{n}$. As noted in proof of Thm 3 in [5], the subgroup $H_{k}\subseteq{\mathfrak{S}}_{n}$ fixing $L_{k}$ is isomorphic to $\mathfrak{S}_{n-k}$. From Lemma 3 in [5], $N_{{\mathfrak{S}}_{n}}(H_{k})\cong\mathfrak{S}_{n-k}\times\mathfrak{S}_{k}$ where $\mathfrak{S}_{k}$ permutes the $k$ roots $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ and $\mathfrak{S}_{n-k}$ permutes the other $n-k$ roots. Now $N_{k}=\tilde{L_{n}}^{N_{{\mathfrak{S}}_{n}}(H_{k})}$. Hence ${\rm Gal}(L_{k}/N_{k})\cong\mathfrak{S}_{k}$ and $[L_{k}:N_{k}]=k!$ and $[N_{k}:K]=\ ^{n}C_{k}$.
Now $K(t_{1},t_{2},\dots,t_{k})\subseteq N_{k}$. Also $L_{k}$ is splitting field of polynomial $x^{k}-t_{1}x^{k-1}+\dots+(-1)^{k}t_{k}$ over $K(t_{1},t_{2},\dots,t_{k})$. Hence $[L_{k}:K(t_{1},t_{2},\dots,t_{k})]\leq k!$. Hence $K(t_{1},t_{2},\dots,t_{k})=N_{k}$.
(2) We give a proof for case (a) only. The proof in the other case follows similarly. Now for any $l<n-1$, we have $\alpha_{i_{l+1}}\not\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{l}})$ and we also have $\alpha_{i_{n}}\in K(\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{n-1}})$ for distinct $i_{m}\leq n$. Since $t_{1}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}$. Hence $t_{1}$ has at most $\ {}^{n}C_{k}$ conjugates inside $\tilde{L_{k}}$ of the form $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}$ for distinct $i_{m}\leq n$.
Hence $[K(t_{1}):K]=\ ^{n}C_{k}$. Now $K(t_{1})\subset N_{k}$. Hence $K(t_{1})=N_{k}$.
(3) For $k<n/2$, we claim $r_{K}(N_{k})=1$. Suppose $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\in K(t_{1})$ for at least one $i_{j}\geq k+1$. By similar argument as in Case 1 of (2), we get a contradiction. Hence the unique descending chain is $L_{k}\supset N_{k}$.
(4) Suppose $k=n/2$ and characteristic of $K$ $\neq 2$. Clearly
Suppose any other $\alpha_{i_{1}}+\alpha_{i_{2}}+\dots+\alpha_{i_{k}}\in K(t_{1})$. Thus at least one $i_{l}$ satisfies $1\leq i_{l}\leq k$. Hence we get that at least one of the $\alpha_{i}$ lies in field generated over $K$ by other $\leq n-2$ many $\alpha_{j}$, which is a contradiction. Hence $r_{K}(N_{k})=2$.
@@ 7.2. Unique ascending chains
We mention an analogue of Thm. 7.1.1 in this subsection. The proof is straightforward.
@@
Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$ be as earlier. Let $t_{K}(L):=[G:H^{G}]$.
@@
    $r_{K}(L)~{}|~{}t_{K}(L)r_{F}(L)$.
@@
    $K=F_{0}\subsetneq F_{1}\subsetneq F_{2}\subset\dots\subsetneq F_{k}$
@@
    $L/K$ is Galois $\iff H^{G}=1\iff H^{G}=H\iff$ the unique descending chain is $K\subsetneq L$.

### v3

@@ 7.1. Unique descending chains
    This unique strictly descending chain of subextensions corresponds to the unique strictly ascending chains of subgroups of $G$
[…]
    such that $N_{G}(H_{k})=H_{k}$. Hence $r_{K}(N_{i})=[H_{i+1}:H_{i}]$ and $s_{K}(N_{i})=[G:H_{i+1}]$.
@@
Proof of (1): Let $N=\tilde{L}^{N_{G}(H)}$. Thus $L/N$ is Galois as $H\unlhd N_{G}(H)$. We have $[L:N]=[N_{G}(H):H]$. Hence by Thm. 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $H\unlhd{\rm Gal}(\tilde{L}/N^{\prime})\subseteq G$. Thus ${\rm Gal}(\tilde{L}/N^{\prime})\subseteq N_{G}(H)$. Hence $N\subseteq N^{\prime}$. Since $[L:N]=[L:N^{\prime}]$, we have $N=N^{\prime}$.
Proof of (2): Let $N_{0}=L$, From (1), we can inductively choose $N_{i}$ for each $i\geq 1$. Let $N_{i}/K$ be the unique intermediate extension of $N_{i-1}/K$ such that $N_{i-1}/N_{i}$ is Galois with degree $[N_{i-1}:N_{i}]=r_{K}(N_{i-1})$. The chain terminates since $L/K$ is finite.
@@
Alternate proof for Thm. 7.1.1 (1).
Let $N=L^{{\rm Aut}(L/K)}$. Hence $L/N$ is Galois and $[L:N]=|{\rm Aut}(L/K)|$. Hence by Thm. 2.1.1 (3), $[L:N]=r_{K}(L)$. Suppose $N^{\prime}/K$ is another intermediate extension such that $L/N^{\prime}$ is Galois and $[L:N^{\prime}]=r_{K}(L)$. Since $L/N^{\prime}$ is Galois, we have $N^{\prime}=L^{{\rm Aut}(L/N^{\prime})}$ and $[L:N^{\prime}]=|{\rm Aut}(L/N^{\prime})|$. Hence $|{\rm Aut}(L/N^{\prime})|=|{\rm Aut}(L/K)|$. By Prop. 6.1.1 (1), ${\rm{\rm Aut}}(L/N^{\prime})\subset{\rm Aut}(L/K)$. Thus ${\rm{\rm Aut}}(L/N^{\prime})={\rm Aut}(L/K)$. Hence $N^{\prime}=N$.
@@
Equivalently we can state Thm. 7.1.1 (1) as follows: There exists a unique intermediate extension $N/K$ such that $L/N$ is Galois of maximum possible degree. This is because, since $L/N$ is Galois, we have $[L:N]=|{\rm Aut}(L/N)|$ and we also have ${\rm{\rm Aut}}(L/N)\subset{\rm Aut}(L/K)$. Hence $[L:N]$ is bounded by $r_{K}(L)$.
@@
Proposition 7.1.4.
Let $L/K$ be a nontrivial finite extension and $N$ be the unique intermediate extension for $L/K$ as in Thm. 7.1.1 (1). Suppose $L=K(\alpha)$ for a primitive element $\alpha\in\bar{K}$ with minimal polynomial $f$ over K such that $r_{K}(f)=r$ and let $\alpha=\alpha_{1},\alpha_{2},\dots,\alpha_{r}$ be the roots of $f$ contained in $L$. Then $N=K(t_{1},t_{2},\ldots,t_{r})$ where $t_{i}$ are elementary symmetric sums of $\alpha_{i}$ for $1\leq i\leq r$.
@@
Let $h$ be the minimal polynomial of $\alpha$ over $N$. The degree of $h$ is $r=|Aut(L/N)|$. Since $Aut(L/N)=Aut(L/K)$, it follows that $h=\Pi_{1\leq i\leq r}(x-\alpha_{i})=x^{r}-t_{1}x^{r-1}+\dots+(-1)^{r}t_{r}$.
Now $K(t_{1},t_{2},\ldots,t_{r})\subset N$. Since $h$ is a polynomial over $K(t_{1},t_{2},\ldots,t_{r})$ which $\alpha$ satisfies. we conclude that $h$ is the minimal polynomial of $\alpha$ over $K(t_{1},t_{2},\ldots,t_{r})$. Thus $N=K(t_{1},t_{2},\ldots,t_{r})$. ∎
@@ 7.2. Unique ascending chains
We mention an analogue of Thm. 7.1.1 in this subsection. The proof is similar.
@@
Let $L/K$ be a nontrivial finite extension. Let $G={\rm Gal}(\tilde{L}/K)$ and $H={\rm Gal}(\tilde{L}/L)$ be as earlier.
@@
    We define the ascending index $t_{K}(L)$ of $L/K$ by $t_{K}(L):=[F:K]$. Let $u_{K}(L):=[L:F]$. Thus $t_{K}(L)\ u_{K}(L)=[L:K]$. We have $t_{K}(L)=[G:H^{G}]$ and $u_{K}(L)=[H^{G}:H]$. We also have $r_{K}(L)~{}|~{}t_{K}(L)\ r_{F}(L)$.
@@
    $K=F_{0}\subsetneq F_{1}\subsetneq F_{2}\subsetneq\dots\subsetneq F_{k}$
@@
    This unique strictly ascending chain of subextensions corresponds to the unique strictly descending chains of subgroups of $G$
[…]
    such that $H^{G_{k}}=G_{k}$. Hence $t_{F_{i}}(L)=[G_{i}:G_{i+1}]$ and $u_{F_{i}}(L)=[G_{i+1}:H]$.
@@
    $L/K$ is Galois $\iff H^{G}=1\iff H^{G}=H\iff$ the unique ascending chain is $K\subsetneq L$.